OpenStudy (anonymous):

solve for C AB=? AC=5 CB=3 ∠C = 90º

OpenStudy (anonymous):

\[\sqrt{32}\]

OpenStudy (anonymous):



OpenStudy (jhannybean):

Use the pythagorean theorem: \(c^2 = a^2+b^2\)

OpenStudy (mathmate):